Long Title

First Author Second Author and Third Author

Abstract

Our premise is ...

Index Terms

keyword 1, keyword 2 ...

I. Introduction

USE the PARstart macro only for the first paragraph in the entire paper...

Theorem 1 (Theorem name) Consider the system ...

Proof: The proof is trivial.

Gregory L. Plett A bio with no face...

filename (missing from disk!)

A. Author is with the Department of Electrical Engineering, Some University, Somewhere CA, 90210, USA, Phone: (xxx) xxx-xxxx, email: xxx@xxxx.xxx.xxx

B. Author is with the Department of Electrical Engineering...

...

December 12, 2005—6 : 50 pm DRAFT